# Learn2Mix: Training Neural Networks Using Adaptive Data Integration

### v2

@@ 4.1 Classification Tasks
Table 2: Test mean squared error (MSE) for learn2mix (L2M) and classical (CL) training.
@@
Dataset MSE (L2M) MSE (CL) MSE (L2M) MSE (CL) MSE (L2M) MSE (CL)
Mean Estimation $1.81\scriptstyle\pm 0.84$ $6.51\scriptstyle\pm 1.52$ $1.45\scriptstyle\pm 0.26$ $1.52\scriptstyle\pm 0.27$ $1.07\scriptstyle\pm 0.09$ $1.17\scriptstyle\pm 0.06$
@@
California Housing $2.52\scriptstyle\pm 0.68$ $2.95\scriptstyle\pm 0.67$ $1.33\scriptstyle\pm 0.32$ $1.82\scriptstyle\pm 0.39$ $0.77\scriptstyle\pm 0.08$ $0.99\scriptstyle\pm 0.10$
@@
Fashion MNIST $89.3\scriptstyle\pm 2.63$ $91.9\scriptstyle\pm 2.37$ $65.1\scriptstyle\pm 1.21$ $70.9\scriptstyle\pm 1.28$ $45.5\scriptstyle\pm 1.21$ $51.6\scriptstyle\pm 1.60$
@@
Figure 3: Comparing model performance errors across six datasets (Mean Estimation, Wine Quality, California Housing, MNIST, Fashion MNIST, CIFAR-10) using MSE Loss for classical training and learn2mix training. The x-axis is the number of elapsed training epochs, while the y-axis is the mean squared error (MSE).
4.2 Regression Tasks
As illustrated in Table 2 and Figure 3, we observe that learn2mix maintains accelerated convergence in the regression context, wherein all the considered datasets are class imbalanced. We first consider the synthetic Mean Estimation dataset, which comprises sets of samples gathered from $k=4$ unique distributions and their associated means. Using the Adam optimizer and Mean Squared Error (MSE) Loss, we train a fully connected network for $E=500$ epochs on Mean Estimation via learn2mix and classical training. We see that the learn2mix-trained neural network observes rapid convergence, achieving a test error below $2.0$ after 100 epochs, at which point the classically-trained network has a test error of $13.0$. For the Wine Quality dataset, we modify the training dataset such that the white wine class has $10\%$ of its original samples. Utilizing the Adam optimizer and MSE Loss, we train a fully connected network for $E=300$ epochs on Wine Quality via learn2mix and classical training. We observe that the learn2mix-trained neural network yields faster convergence, achieving a test error below $2.5$ after 200 epochs, at which point the classically-trained network has a test error of $5.0$. Finally, on the California Housing dataset, we modify the training dataset such that three of the classes have $5\%$ of their original samples. Using the Adam optimizer and MSE Loss, we train a fully connected network for $E=1200$ epochs on California Housing via learn2mix and classical training. We again see that the learn2mix-trained network converges faster, achieving a test error below $0.8$ after 1200 epochs, while the classically-trained network has a test error of $0.99$. These empirical evaluations support our previous intuition pertaining to the extension of learn2mix to imbalanced regression settings.
4.3 Image Reconstruction Tasks
Per Table 2 and Figure 3, we note that the class-imbalanced image reconstruction tasks also observe faster convergence using learn2mix. For the MNIST case, we modify the training dataset such that half of the classes retain $20\%$ of their original samples. Leveraging the Adam optimizer and MSE Loss, we train an autoencoder for $E=40$ epochs on MNIST using learn2mix and classical training. We observe that the learn2mix-trained autoencoder exhibits improved convergence, achieving a test error less than $1.0$ after 35 epochs, which the classically-trained autoencoder achieves after 40 epochs. Correspondingly, for Fashion MNIST, we modify the training dataset such that half of the classes retain $20\%$ of their original samples (paralleling MNIST). Using the Adam optimizer and MSE Loss, we train an autoencoder for $E=70$ epochs on Fashion MNIST, leveraging learn2mix and classical training. We observe that the learn2mix-trained autoencoder converges faster, achieving a test error below $54.0$ after 50 epochs, which the classically-trained autoencoder achieves after 65 epochs. We also consider CIFAR-10, wherein we modify the training dataset such that all but two classes retain $20\%$ of their original samples. Utilizing the Adam optimizer and MSE Loss, we train an autoencoder for $E=110$ epochs on CIFAR-10, leveraging learn2mix and classical training. We observe that the learn2mix-trained autoencoder also converges faster and achieves a test error below $148.0$ after 100 epochs, which the classically-trained autoencoder achieves after 110 epochs. Cumulatively, these empirical evaluations demonstrate the improved performance yielded by learn2mix trained models over classically trained models in constrained training regimes.
@@ 5 Conclusion
In this work, we presented learn2mix, a new training strategy that adaptively modifies class proportions in batches via real-time class-wise error rates, accelerating model convergence. We formalized the learn2mix mechanism through a bilevel optimization framework, and outlined its theoretical advantages in aligning class proportions with optimal error rates. Empirical evaluations across classification, regression, and reconstruction tasks on both balanced and imbalanced datasets confirmed that learn2mix not only accelerates convergence compared to classical training methods, but also reduces overfitting in the presence of class-imbalances. As a consequence, models trained with learn2mix achieved improved performance in constrained training regimes. Our findings underscore the potential of dynamic batch composition strategies in optimizing neural network training, paving the way for more efficient and robust machine learning models in resource-constrained environments.
Acknowledgements.
This work was supported in part by the U.S. Air Force Office of Scientific Research (AFOSR) under award FA8750-20-2-0504. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the U.S. Department of Defense.
@@ Appendix A Proofs of the Theoretical Results
$\theta^{t+1}=\theta^{t}-\eta\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t}),\quad\quad\text{with: \quad$\eta\in\mathbb{R}_{>0}$},$ (14)
@@
Taking the limit and re-indexing the summation using $n=t-1-l$ and $l=t-1-n$, we obtain:
@@
[…]
@@
[…]
@@
Consequently, suppose that $\mathcal{H}(\theta^{t},\theta^{*},\eta,\tilde{\alpha},\alpha^{t})=2\eta\big{[}(\theta^{t}-\theta^{*})^{T}(\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})-\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha}))\big{]}$, and let $\mathcal{J}(\theta^{t},\eta,\tilde{\alpha},\alpha^{t})=\eta^{2}\big{[}\|\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})\|^{2}-\|\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha})\|^{2}\big{]}$. Suppose the loss gradient, $\smash{\nabla_{\theta}\mathcal{L}(\theta,\alpha)}$, is Lipschitz continuous in $\alpha$ with Lipschitz constant, $L_{\alpha}$. We now upper bound $\mathcal{J}(\theta^{t},\eta,\tilde{\alpha},\alpha^{t})$:
$\mathcal{J}(\theta^{t},\eta,\alpha,\alpha^{t})$ $=\eta^{2}\big{[}\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})-\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha})\big{]}^{T}\big{[}\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})+\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha})\big{]}$
$\leq\|\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})-\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha})\|\|\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})+\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha})\|$ (44)
$\leq 2\eta^{2}L_{\alpha}\|\alpha^{t}-\tilde{\alpha}\|\Big{[}\|\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})\|+\|\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha})\|\Big{]}$ (45)
@@
$\mathcal{H}(\theta^{t},\theta^{*},\eta,\tilde{\alpha},\alpha^{t})$ $=2\eta\Big{[}(\theta^{t}-\theta^{*})^{T}\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})-(\theta^{t}-\theta^{*})^{T}\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha})\Big{]}$ (49)
$\geq 2\eta\Big{[}(\theta^{t}-\theta^{*})^{T}\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})-\|\theta^{t}-\theta^{*}\|\|\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha})\|\Big{]}$ (50)
@@
[…]
@@
$\begin{split}\mathcal{K}(\theta^{t},\theta^{*},\eta,\gamma,\tilde{\alpha},\alpha^{t})\leq 0&\implies\big{[}\mathcal{F}(\theta^{t},\theta^{*},\eta,\alpha^{t})\big{]}^{2}-\big{[}\mathcal{G}(\theta^{t},\theta^{*},\eta,\tilde{\alpha})\big{]}^{2}\leq 0\\ &\implies\big{[}\mathcal{F}(\theta^{t},\theta^{*},\eta,\alpha^{t})\big{]}\leq\big{[}\mathcal{G}(\theta^{t},\theta^{*},\eta,\tilde{\alpha})\big{]}.\end{split}$ (56)
@@
[…]
@@
Appendix B Dataset Descriptions
B.1 MNIST Dataset
@@
The MNIST (Modified National Institute of Standards and Technology) dataset is a collection of handwritten digits commonly used to train image processing systems. For the MNIST classification result from Section 4.1, the original training dataset, $J$, comprises $N=60000$ samples, wherein the fixed-proportion mixing parameters (for default numerical class ordering of digits from $1-10$) are:
@@
The test dataset, $K$, comprises $N_{\text{test}}=10000$ samples, with class proportions equivalent to the class proportions in the base MNIST test dataset. For MNIST reconstruction (see Section 4.3), we utilize manual class imbalancing, reducing the number of samples comprising each numerical class $6-10$ by a factor of 5. The original training dataset, $J$, now contains $N=36475$ samples, wherein the fixed-proportion mixing parameters (for default numerical class ordering of digits from $1-10$) are:
@@
B.2 Fashion MNIST Dataset
The Fashion MNIST dataset is a collection of clothing images commonly used to train image processing systems. For the Fashion MNIST classification result from Section 4.1, the original training dataset, $J$, consists of $N=60000$ samples, wherein the fixed-proportion mixing parameters (for default numerical class ordering of clothing from $1-10$) are:
@@ B.2 Fashion MNIST Dataset
The test dataset, $K$, comprises $N_{\text{test}}=10000$ samples, with class proportions equivalent to the class proportions in the base Fashion MNIST test dataset. For Fashion MNIST reconstruction (see Section 4.3), we use manual class imbalancing, reducing the number of samples within each numerical class $6-10$ by a factor of 5. The original training dataset $J$, now has $N=36000$ samples. The fixed-proportion mixing parameters (for default numerical class ordering of clothing from $1-10$) are:
@@
B.3 CIFAR-10 Dataset
The CIFAR-10 dataset is a collection of color images categorized into $10$ different classes, and is commonly used to train image processing systems. For the CIFAR-10 classification result in Section 4.1, the original training dataset, $J$, comprises $N=50000$ samples, wherein the fixed-proportion mixing parameters (for default numerical class ordering of categories from $1-10$) are:

### v3

@@ 4.3 Image Reconstruction Tasks
Table 2: Test mean squared error for learn2mix (L2M) and classical (CL) training.
@@
Dataset Err (L2M) Err (CL) Err (L2M) Err (CL) Err (L2M) Err (CL)
Mean Estim. $1.81\scriptstyle\pm 0.84$ $6.51\scriptstyle\pm 1.52$ $1.45\scriptstyle\pm 0.26$ $1.52\scriptstyle\pm 0.27$ $07\scriptstyle\pm 0.09 1.1$$17\scriptstyle\pm 0.06$
@@
Cali. Housing $2.52\scriptstyle\pm 0.68$ $2.95\scriptstyle\pm 0.67$ $1.33\scriptstyle\pm 0.32$ $1.82\scriptstyle\pm 0.39$ $0.77\scriptstyle\pm 0.08$ $0.99\scriptstyle\pm 0.10$
@@
Fsh. MNIST $89.3\scriptstyle\pm 2.63$ $91.9\scriptstyle\pm 2.37$ $65.1\scriptstyle\pm 1.21$ $70.9\scriptstyle\pm 1.28$ $45.5\scriptstyle\pm 1.21$ $51.6\scriptstyle\pm 1.60$
@@
Figure 3: Comparing model performance errors for classical training and learn2mix training. The x-axis is the number of elapsed training epochs, while the y-axis is the mean squared error (MSE).
@@ 5 Conclusion
In this work, we presented learn2mix, a new training strategy that adaptively modifies class proportions in batches via real-time class-wise error rates, accelerating model convergence. We formalized the learn2mix mechanism through a bilevel optimization framework, and outlined its theoretical advantages in aligning class proportions with optimal error rates. Empirical evaluations across classification, regression, and reconstruction tasks on both balanced and imbalanced datasets confirmed that learn2mix not only accelerates convergence compared to classical training methods, but also reduces overfitting in the presence of class-imbalances. Accordingly, models trained with learn2mix achieved improved performance in constrained training regimes. Our findings underscore the potential of dynamic batch composition strategies in optimizing neural network training, paving the way for more efficient and robust machine learning models in resource-constrained environments.
Acknowledgments and Disclosure of Funding
Shyam Venkatasubramanian and Vahid Tarokh were supported in part by the Air Force Office of Scientific Research under award FA9550-21-1-0235.
@@ Appendix A Proofs of the Theoretical Results
$\theta^{t+1}=\theta^{t}-\eta\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t}),\quad\quad\text{with: \qquad$\eta\in\mathbb{R}_{>0}$},$ (14)
@@
Taking the limit and re-indexing the summation using n=t1ln=t-1-l and l=t1nl=t-1-n, we obtain:
@@
$\text{Wherein: }\|\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})\|+\|\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha})\|\leq 2L^{*}\|\theta^{t}-\theta^{*}\|.$ (29)
@@
$\text{Where: }\beta=\frac{\big(\frac{\mu^{*}}{2}-L^{*}\big)\|\theta^{t}-\theta^{*}\|^{2}+\tilde{\alpha}^{T}(\mathcal{L}(\theta^{t})-\mathcal{L}(\theta^{*}))}{\eta L_{\alpha}L^{*}\Big\|\frac{\mathcal{L}(\theta^{t-1})}{\mathds{1}_{k}^{T}\mathcal{L}(\theta^{t-1})}-\tilde{\alpha}\Big\|\Big[\|\theta^{t}-\theta^{*}\|-(\mathcal{L}(\theta^{t})-\mathcal{L}(\theta^{*}))\Big]}$ (40)
@@
Consequently, suppose that $)]\mathcal{H}(\theta^{t},\theta^{*},\eta,\tilde{\alpha},\alpha^{t})=2\eta\big[(\theta^{t}-\theta^{*})^{T}(\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})-\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha}$$)\big$], and let $\mathcal{J}(\theta^{t},\eta,\tilde{\alpha},\alpha^{t})=\eta^{2}\big[\|\nabla_{\theta}\mathcal{L}(\theta^{t},\alpha^{t})\|^{2}-\|\nabla_{\theta}\mathcal{L}(\theta^{t},\tilde{\alpha})\|^{2}\big]$. Suppose the loss gradient, $\smash{\nabla_{\theta}\mathcal{L}(\theta,\alpha)}$, is Lipschitz continuous in $\alpha$ with Lipschitz constant, $L_{\alpha}$. We now upper bound $\mathcal{J}(\theta^{t},\eta,\tilde{\alpha},\alpha^{t})$:
[…]
@@
[…]
@@
$\text{Where: }\mathcal{K}(\theta^{t},\theta^{*},\eta,\gamma,\tilde{\alpha},\alpha^{t})=-2\eta\bigg[\Big(\frac{\mu^{*}}{2}-L^{*}\Big)\|\theta^{t}-\theta^{*}\|^{2}+\tilde{\alpha}^{T}(\mathcal{L}(\theta^{t})-\mathcal{L}(\theta^{*}))$
@@
[…]
@@
$\text{Where: }\beta=\frac{\big(\frac{\mu^{*}}{2}-L^{*}\big)\|\theta^{t}-\theta^{*}\|^{2}+\tilde{\alpha}^{T}(\mathcal{L}(\theta^{t})-\mathcal{L}(\theta^{*}))}{\eta L_{\alpha}L^{*}\Big\|\frac{\mathcal{L}(\theta^{t-1})}{\mathds{1}_{k}^{T}\mathcal{L}(\theta^{t-1})}-\tilde{\alpha}\Big\|\Big[\|\theta^{t}-\theta^{*}\|-(\mathcal{L}(\theta^{t})-\mathcal{L}(\theta^{*}))\Big]}.$ (59)
@@
Appendix B Additional Empirical Results
For further performance verification of learn2mix, we present several ablation studies quantifying the effects of different architectures, optimizers, batch sizes, learning rates, and mixing rates for the considered classification tasks from the main text. We further present the worst-class classification accuracy on Imagenette to further gauge the efficacy of learn2mix within imbalanced classification settings, and illustrate how the mixing parameters converge to a stable distribution on Mean Estimation. We first consider CIFAR-10 and CIFAR-100 (per Section C), and evaluate whether the gains afforded by learn2mix persist across architectures. For CIFAR-10, we recall the Large LeNet architecture, trained using the Adam optimizer and Cross Entropy Loss with learning rate $\eta=7\text{e-}5$ for $\underline{E}=200$ s, and the MobileNet-V3 Small architecture, trained using the Adam optimizer and Cross Entropy Loss with learning rate $\eta=1\text{e-}4$ for $\underline{E}=750$ s. For CIFAR-100, we consider again the MobileNet-V3 Small architecture, trained using the Adam optimizer and Cross Entropy Loss with learning rate $\eta=1\text{e-}4$ for $\underline{E}=200$ s, and the Large LeNet architecture, trained using the Adam optimizer and Cross Entropy Loss with learning rate $\eta=1\text{e-}4$ for $\underline{E}=50$ s. The results are depicted in Figure 4. We observe that for both Large LeNet and MobileNet-V3 Small, the learn2mix-trained models converge faster than the classical, FCL, SMOTE, IS, and CURR trained models.
(a) CIFAR-10: Large LeNet
(b) CIFAR-10: MobileNet-V3 Small
(c) CIFAR-100: Large LeNet
(d) CIFAR-100: MobileNet-V3 Small
Figure 4: Comparing model classification errors for learn2mix, classical, FCL, SMOTE, IS, and CURR training. The x-axis is the elapsed [training] time, while the y-axis is the classification error.
Next, we evaluate the robustness of learn2Mix to different optimizers and batch sizes. As we used the Adam optimizer in the main text, we now consider the RMSProp optimizer [Graves, 2013] with batch size $M\in\{250,500,1000\}$. We train LeNet-5 on MNIST using Cross Entropy Loss with learning rate $\eta=1\text{e-}5$ for $\underline{E}=45$ s, $\underline{E}=60$ s, and $\underline{E}=70$ s. As depicted in Figure 5, we see that learn2mix converges faster than classical, FCL, SMOTE, IS, and CURR training.
(a) MNIST: $M=150$
(b) MNIST: $M=500$
(c) MNIST: $M=1000$
Figure 5: Comparing model classification errors for learn2mix, classical, FCL, SMOTE, IS, and CURR training. The x-axis is the elapsed [training] time, while the y-axis is the classification error.
We further verify the robustness of learn2Mix to different learning rates. We train LeNet-5 on MNIST using Cross Entropy Loss with learning rate $\eta\in\{1\text{e-}5,1\text{e-}4,1\text{e-}3\}$ for $\underline{E}=75$ s, $\underline{E}=50$ s, and $\underline{E}=45$ s. Per Figure 6, we observe that the faster convergence afforded by learn2mix is apparent for $\eta\in\{1\text{e-}5,1\text{e-}4\}$. For $\eta=1\text{e-}3$, we note that after convergence, the learn2mix train error continues to decreases at a faster rate than the the classical, FCL, SMOTE, IS, and CURR train errors.
(a) MNIST: $\eta=1\text{e-}5$
(b) MNIST: $\eta=1\text{e-}4$
(c) MNIST: $\eta=1\text{e-}3$
Figure 6: Comparing model classification errors for learn2mix, classical, FCL, SMOTE, IS, and CURR training. The x-axis is the elapsed [training] time, while the y-axis is the classification error.
We now illustrate the worst-class classification accuracy on Imagenette and IMDB as an additional metric to gauge the efficacy of learn2mix for imbalanced classification settings. We train ResNet-18 on Imagenette via Cross Entropy Loss with learning rate $\eta=1\text{e-}5$ for $\underline{E}=240$ s, and a transformer on IMDB using Cross-Entropy Loss with learning rate $\eta=1\text{e-}4$ for $\underline{E}=150$ s, and record the test classification accuracy of the worst class after each training epoch, $t$. To demonstrate relative insensitivity to the choice of $\gamma$, we ablate the mixing rate for $\gamma\in[0.01,0.1]$. The result is depicted in Figure 7. We see that learn2mix offers a considerable improvement in the worst-class classification accuracy metric versus classical, FCL, SMOTE, IS, and CURR training, which matches intuition; the theoretical foundation of learn2Mix is to increase the proportion of harder classes during training, which directly translates to stronger results for the most challenging classes.
(a) Imagenette
(b) IMDB
Figure 7: Comparing worst-class model classification accuracies using learn2mix, classical, FCL, SMOTE, IS, and CURR training on Imagenette and IMDB. The x-axis is the elapsed [training] time, while the y-axis is the classification accuracy of the worst-class.
To illustrate how the mixing parameters converge to a stable distribution during training (as detailed in Section 2), we train a fully connected network on Mean Estimation (where the Normal, Exponential, and Chi-squared cases have similar variance but the Uniform case is substantially more variable) using Cross-Entropy Loss with learning rate $\eta=5\text{e-}5$ for $E=500$ epochs. As depicted in Figure 8, learn2mix prioritizes the hardest class without overstating differences among the easier ones.
Figure 8: Evolution of learn2mix mixing parameters across training epochs on Mean Estimation.
Appendix C Dataset Descriptions
C.1 MNIST Dataset
@@
The MNIST (Modified National Institute of Standards and Technology) dataset is a collection of handwritten digits commonly used to train image processing systems. For the MNIST classification result from Section 4.1, the original training dataset, $J$, comprises $N=60000$ samples, wherein the fixed-proportion mixing parameters (for default numerical class ordering of digits from 1101-10) are:
@@
The test dataset, $K$, comprises $N_{\text{test}}=10000$ samples, with class proportions equivalent to the class proportions in the base MNIST test dataset. For MNIST reconstruction (see Section 4.3), we utilize manual class imbalancing, reducing the number of samples comprising each numerical class 6106-10 by a factor of 5. The original training dataset, $J$, now contains $N=36475$ samples, wherein the fixed-proportion mixing parameters (for default numerical class ordering of digits from 1101-10) are:
@@
C.2 Fashion MNIST Dataset
The Fashion MNIST dataset is a collection of clothing images commonly used to train image processing systems. For the Fashion MNIST classification result from Section 4.1, the original training dataset, $J$, consists of $N=60000$ samples, wherein the fixed-proportion mixing parameters (for default numerical class ordering of clothing from 1101-10) are:
@@ C.2 Fashion MNIST Dataset
The test dataset, $K$, comprises $N_{\text{test}}=10000$ samples, with class proportions equivalent to the class proportions in the base Fashion MNIST test dataset. For Fashion MNIST reconstruction (see Section 4.3), we use manual class imbalancing, reducing the number of samples within each numerical class 6106-10 by a factor of 5. The original training dataset $J$, now has $N=36000$ samples. The fixed-proportion mixing parameters (for default numerical class ordering of clothing from 1101-10) are:
@@
C.3 CIFAR-10 Dataset
The CIFAR-10 dataset is a collection of color images categorized into $10$ different classes, and is commonly used to train image processing systems. For the CIFAR-10 classification result in Section 4.1, the original training dataset, $J$, comprises $N=50000$ samples, wherein the fixed-proportion mixing parameters (for default numerical class ordering of categories from 1101-10) are: